Fungicides account for 1/10 of the perticides used in the United States. Insecticides account for 1/4 of all the pesticides in the United States. The ratio of fungicides to insecticides used in the United States can be written 1/10 / 1/4. Write this ratio in simplest form.
1/10 / 1/4
First we flip the last problem:
1/10 / 4/1
We then cross out the 10 and 4 by dividing 2 in to it leaving us with:
1/5 / 2/1
From here I am not sure where to go unless I have already messed up. Hope you can help me out. Thanks a bunch

Answer by stanbon(75887)

1/10 / 1/4
First we flip the last problem:
Comment: You invert the denominator and MULTIPLY
=1/10 * 4/1
We then cross out the 10 and 4 by dividing 2 in to it leaving us with:
=1/5 * 2/1
=2/5
Cheers,
Stan H.